# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ Appendix C Proof of Theorem 3
To simplify notation, let $\tilde{p}(x_{1},\cdots,x_{d})$ denote the distribution of the extended sequence $(X_{1},\cdots,X_{d})$. For $\mathbf{s}\in\mathcal{X}^{d}$, we let $h_{\mathbf{s}}(x_{1},\cdots,x_{k})$ denote the distribution of $k$ draws without replacement from an urn containing the elements of $\mathbf{s}$ and let $m_{\mathbf{s}}(x_{1},\cdots,x_{k})$ denote the distribution of $k$ draws with replacement from the same urn. Note that $m_{\mathbf{s}}(x_{1},\cdots,x_{k})$ is equivalent to the empirical distribution of $\mathbf{s}$.
Using the extended distribution $\tilde{p}$, the source distribution is equivalent to the following weighted sum of urn distributions:
@@
Using the log-sum inequality, the relative entropy can be upper bounded as
$\begin{split}&D(p\|q)=\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})\log\left(\frac{p(\mathbf{x})}{q(\mathbf{x})}\right)\\ &=\sum_{\mathbf{x}\in\mathcal{X}^{k}}\left[\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})h_{\mathbf{s}}(\mathbf{x})\right]\log\left(\frac{\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})h_{\mathbf{s}}(\mathbf{x})}{\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})m_{\mathbf{s}}(\mathbf{x})}\right)\\ &\leq\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})\sum_{\mathbf{x}\in\mathcal{X}^{k}}h_{\mathbf{s}}(\mathbf{x})\log\left(\frac{h_{\mathbf{s}}(\mathbf{x})}{m_{\mathbf{s}}(\mathbf{x})}\right)\\ &=\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})D(h_{\mathbf{s}}\|m_{\mathbf{s}}),\end{split}$ (80)
From Theorem $4.5$ of [8], we know that for all $\mathbf{s}\in\mathcal{X}^{d}$,
@@
Therefore $D(p\|q)$ is bounded above by the same quantity.
Next, we apply a similar line of reasoning as the proof of Theorem 2 by considering collision-free sampling patterns. Let $\mathbf{W}=(W_{1},\cdots,W_{k})$ be i.i.d. uniformly over $[d]$ and let $\mathbf{Y}=(X_{W_{1}},\cdots,X_{W_{k}})$. By definition, $\mathbf{Y}$ is distributed according to $q(\mathbf{x})$. Next, define the set of all collision-free sample patterns as =$\mathcal{F}=\{\mathbf{w}\in[d]^{k}\ |\ w_{i}\neq w_{j}\text{ for all }i\neq j\}$. It follows that:
$q(\mathbf{x})$ $=\sum_{\mathbf{w}\in[d]^{k}}\mathrm{Pr}(\mathbf{W}=\mathbf{w})\mathrm{Pr}(\mathbf{Y}=\mathbf{x}|\mathbf{W}=\mathbf{w})$
$\geq\sum_{\mathbf{w}\in\mathcal{F}}\mathrm{Pr}(\mathbf{W}=\mathbf{w})\mathrm{Pr}(\mathbf{Y}=\mathbf{x}|\mathbf{W}=\mathbf{w})$
$=\sum_{\mathbf{w}\in\mathcal{F}}\frac{1}{d^{k}}p(\mathbf{x})=\frac{d^{\underline{k}}}{d^{k}}p(\mathbf{x}).$ (82)
@@ Appendix D Converse for Scheduling
We prove (33) using a volume bound argument on the total number of possible scheduling patterns in relation to the maximum number of scheduling patterns can correspond to a single codeword. This allows us to show that the optimal common message rate is bounded from below as (33), i.e.,
@@
$\begin{split}\mathbf{v}^{*}=\operatorname*{arg\,max}_{\mathbf{v}}\sum_{\mathcal{U}\in{{[b]}\choose{k}}}\prod_{\ell\in\mathcal{U}}v_{\ell}&=\operatorname*{arg\,max}_{\mathbf{v}}\sum_{\mathcal{U}\in{{[b]}\choose{k}}}k!\prod_{\ell\in\mathcal{U}}\frac{v_{\ell}}{n}.\end{split}$ (87)
@@
$\begin{split}H(T)&=-\sum_{t}\mathrm{Pr}(T=t)\log(\mathrm{Pr}(T=t))\\ &\geq-\sum_{t}\mathrm{Pr}(T=t)\log\left(\frac{n^{\underline{k}}}{b^{k}n^{k}}\right)\\ &=k\log(b)-\log\left(\frac{n^{k}}{n^{\underline{k}}}\right).\end{split}$ (90)
@@ Appendix E Converse for Categorization
We use a volume bound argument to show (37), i.e. for the problem of categorization, the optimal common message rate is bounded from below as
@@
$\begin{split}\operatorname*{arg\,max}_{\mathbf{v}}\prod_{\ell=1}^{c}\frac{v_{\ell}^{k_{\ell}}}{k_{\ell}!}&=\operatorname*{arg\,min}_{\mathbf{v}}-\sum_{\ell=1}^{c}\frac{k_{\ell}}{k}\log\left(\frac{v_{\ell}}{n}\right).\end{split}$ (94)
@@
$\begin{split}H(T)&=-\sum_{t}\mathrm{Pr}(T=t)\log(\mathrm{Pr}(T=t))\\ &\geq-\sum_{t}\mathrm{Pr}(T=t)\log\left(\frac{d_{\text{max}}}{{{n}\choose{k}}{{k}\choose{k_{1}\cdots k_{c}}}}\right)\\ &=\log\left({{n}\choose{k}}{{k}\choose{k_{1}\cdots k_{c}}}\frac{1}{d_{\text{max}}}\right)\\ &\geq kH(\rho)-\log\left(\frac{n^{k}}{n^{\underline{k}}}\right).\end{split}$ (97)
@@ Appendix F Proof of Theorem 4
$f_{\mathbf{m}}(x_{1},\cdots x_{K},a_{1},\cdots a_{K})=$ $\min$ $t$
@@
Let $T=f_{\mathbf{M}}(X_{1},\cdots,X_{K},A_{1},\cdots,A_{K})$ be the index of the first matching codeword for a random source and a random activity pattern, along with a randomly generated codebook $\mathbf{M}$, where the entries of the codewords in $\mathbf{M}$ are generated in an i.i.d. fashion according to $p(x)$. By the same logic as the proof of Theorem 1, we know that
@@ Appendix G Proof of Theorem 6
In order to lower bound the relative entropy $D(h\|q)$, we first lower bound the cross-entropy using Jensen’s inequality, i.e.
$\begin{split}\sum_{\mathbf{x}\in[d]^{k}}h(\mathbf{x})\log\left(\frac{1}{q(\mathbf{x})}\right)&=\mathbb{E}\left[\log\left(\frac{1}{q(\mathbf{x})}\right)\right]\\ &\geq\log\left(\frac{1}{\mathbb{E}[q(\mathbf{x})]}\right).\end{split}$ (102)
@@
=$\mathcal{B}=\{\mathbf{x}\in[d]^{k}\ :\ x_{i}=x_{j}\ \text{for some }i\neq j\}.$ (103)
By definition of $h$, we know that $h(\mathbf{x})=0$ for all $\mathbf{x}\in\mathcal{B}$. Therefore
$\mathbb{E}[q(\mathbf{x})]=\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}h(\mathbf{x})q(\mathbf{x})=\left({{d}\choose{k}}k!\right)^{-1}\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}q(\mathbf{x}).$ (104)
From [5], we know that the i.i.d. mixture distribution that maximizes $\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}q(\mathbf{x})$ is the distribution that is uniform over $[d]^{k}$. Therefore
@@
$\begin{split}D(h\|q)&=\mathbb{E}\left[\log\left(\frac{1}{q(\mathbf{x})}\right)\right]-H(\mathbf{X})\\ &\geq\log\left({{d}\choose{k}}k!\right)+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)-H(\mathbf{X})\\ &=\log\left(\frac{d^{k}}{d^{\underline{k}}}\right).\end{split}$ (106)

### v4

@@ Appendix C Proof of Theorem 3
To simplify notation, let $\tilde{p}(x_{1},\cdots,x_{d})$ denote the distribution of the extended sequence $(X_{1},\cdots,X_{d})$. For $\mathbf{s}\in\mathcal{X}^{d}$, we let $h_{\mathbf{s}}(x_{1},\cdots,x_{k})$ denote the distribution of $k$ draws without replacement from an urn containing the elements of $\mathbf{s}$ and let $m_{\mathbf{s}}(x_{1},\cdots,x_{k})$ denote the distribution of $k$ draws with replacement from the same urn. Note that $m_{\mathbf{s}}(x_{1},\cdots,x_{k})$ is equivalent to the i.i.d. distribution with the empirical distribution of $\mathbf{s}$ as its marginal.
Using the extended distribution $\tilde{p}$, the distribution of the sources is equivalent to the following weighted sum of urn distributions:
@@
Using the log-sum inequality, the KL divergence can be upper bounded as
$D($ $p\|q)=\sum_{\mathbf{x}\in\mathcal{X}^{k}}p(\mathbf{x})\log\left(\frac{p(\mathbf{x})}{q(\mathbf{x})}\right)$ (129)
$=\sum_{\mathbf{x}\in\mathcal{X}^{k}}\left(\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})h_{\mathbf{s}}(\mathbf{x})\right)\log\left(\frac{\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})h_{\mathbf{s}}(\mathbf{x})}{\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})m_{\mathbf{s}}(\mathbf{x})}\right)$ (130)
$\leq\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})\sum_{\mathbf{x}\in\mathcal{X}^{k}}h_{\mathbf{s}}(\mathbf{x})\log\left(\frac{h_{\mathbf{s}}(\mathbf{x})}{m_{\mathbf{s}}(\mathbf{x})}\right)$ (131)
$=\sum_{\mathbf{s}\in\mathcal{X}^{d}}\tilde{p}(\mathbf{s})D(h_{\mathbf{s}}\|m_{\mathbf{s}}).$ (132)
From [10, Theorem 4.5], we know that for all $\mathbf{s}\in\mathcal{X}^{d}$,
@@
Therefore, $D(p\|q)$ is bounded from above by the same quantity. This line of reasoning is similar to the reasoning used in [14] to prove (9).
Next, we apply a similar line of reasoning as the proof of Theorem 2 by considering collision-free sampling patterns. We represent the sampling process by defining the positions of $\mathbf{s}$ that are sampled as $\mathbf{w}=(w_{1},\cdots,w_{k})$ where each $w_{i}$ is i.i.d. over $[d]$. Likewise, we define the set of all collision-free sample patterns as =$\mathcal{F}=\{\mathbf{w}\in[d]^{k}\ |\ w_{i}\neq w_{j}\text{ for all }i\neq j\}$. It follows that
$q(\mathbf{x})$ $=\sum_{\mathbf{s}\in\mathcal{X}^{k}}\sum_{\mathbf{w}\in[d]^{k}}p(\mathbf{s})p(\mathbf{w})\mathbbm{1}({\mathbf{x}=\mathbf{s}_{\mathbf{w}}})$ (134)
$\geq\sum_{\mathbf{s}\in\mathcal{X}^{k}}\sum_{\mathbf{w}\in\mathcal{F}}p(\mathbf{s})p(\mathbf{w})\mathbbm{1}({\mathbf{x}=\mathbf{s}_{\mathbf{w}}})$ (135)
$=\sum_{\mathbf{w}\in\mathcal{F}}\frac{1}{d^{k}}p(\mathbf{x})$ (136)
$=\frac{d^{\underline{k}}}{d^{k}}p(\mathbf{x}).$ (137)
@@ Appendix D Converse for Scheduling
We prove (49) using a volume bound argument on the total number of possible scheduling patterns in relation to the maximum number of scheduling patterns can correspond to a single codeword. This allows us to show that the optimal common message length is bounded from below as (49), i.e.,
@@
$\mathbf{v}^{*}$ $=\operatorname*{arg\,max}_{\mathbf{v}}\sum_{\mathcal{U}\in{{[b]}\choose{k}}}\prod_{\ell\in\mathcal{U}}v_{\ell}$ (142)
$=\operatorname*{arg\,max}_{\mathbf{v}}\sum_{\mathcal{U}\in{{[b]}\choose{k}}}k!\prod_{\ell\in\mathcal{U}}\frac{v_{\ell}}{n}.$ (143)
@@
$H(T)$ $=-\sum_{t}\mathrm{Pr}(T=t)\log(\mathrm{Pr}(T=t))$ (146)
$\geq-\sum_{t}\mathrm{Pr}(T=t)\log\left(\frac{n^{\underline{k}}}{b^{k}n^{k}}\right)$ (147)
$=k\log(b)-\log\left(\frac{n^{k}}{n^{\underline{k}}}\right).$ (148)
@@ Appendix E Converse for Categorization
We use a volume bound argument to show (56), i.e. for the problem of categorization, the optimal common message length is bounded from below as
@@
$\operatorname*{arg\,max}_{\mathbf{v}}\prod_{\ell=1}^{c}\frac{v_{\ell}^{k_{\ell}}}{k_{\ell}!}=\operatorname*{arg\,min}_{\mathbf{v}}-\sum_{\ell=1}^{c}\frac{k_{\ell}}{k}\log\left(\frac{v_{\ell}}{n}\right).$ (152)
@@
$H(T)$ $=-\sum_{t}\mathrm{Pr}(T=t)\log(\mathrm{Pr}(T=t))$ (155)
$\geq-\sum_{t}\mathrm{Pr}(T=t)\log\left(\frac{d_{\text{max}}}{{{n}\choose{k}}{{k}\choose{k_{1}\cdots k_{c}}}}\right)$ (156)
$=\log\left({{n}\choose{k}}{{k}\choose{k_{1}\cdots k_{c}}}\frac{1}{d_{\text{max}}}\right)$ (157)
$\geq kH(\rho)-\log\left(\frac{n^{k}}{n^{\underline{k}}}\right).$ (158)
@@ Appendix F Proof of Theorem 4
$g_{\mathbf{m}}(x_{1},\cdots x_{K},a_{1},\cdots a_{K})=$ $\min$ $t$
@@
Let $T=g_{\mathbf{M}}(X_{1},\cdots,X_{K},A_{1},\cdots,A_{K})$ be the index of the first matching codeword using randomly generated codebook $\mathbf{M}$. Since the sources are distributed i.i.d. according to $p(x)$, we choose to generate the entries of the codewords in $\mathbf{M}$ in an i.i.d. fashion according to $p(x)$ as well. By the same argument as the proof of Theorem 1, we have that
@@ Appendix G Proof of Theorem 6
In order to lower bound the KL divergence $D(r\|q)$, we first lower bound the cross-entropy using Jensen’s inequality, i.e.,
$\sum_{\mathbf{x}\in[d]^{k}}r(\mathbf{x})\log\left(\frac{1}{q(\mathbf{x})}\right)$ $=\mathbb{E}\left[\log\left(\frac{1}{q(\mathbf{x})}\right)\right]$ (164)
$\geq\log\left(\frac{1}{\mathbb{E}[q(\mathbf{x})]}\right).$ (165)
@@
=$\mathcal{B}=\{\mathbf{x}\in[d]^{k}\ |\ x_{i}=x_{j}\ \text{for some }i\neq j\}.$ (166)
By definition of $r$, we know that $r(\mathbf{x})=0$ for all $\mathbf{x}\in\mathcal{B}$. Therefore
$\mathbb{E}[q(\mathbf{x})]=\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}r(\mathbf{x})q(\mathbf{x})=\left({{d}\choose{k}}k!\right)^{-1}\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}q(\mathbf{x}).$ (167)
From [5], we know that the i.i.d. mixture distribution that maximizes $\sum_{\mathbf{x}\in\mathcal{B}^{\sf c}}q(\mathbf{x})$ is the uniform distribution over $[d]^{k}$. Therefore
@@
$D(r\|q)$ $=\mathbb{E}\left[\log\left(\frac{1}{q(\mathbf{x})}\right)\right]-H(\mathbf{X})$ (169)
$\geq\log\left({{d}\choose{k}}k!\right)+\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)-H(\mathbf{X})$ (170)
$=\log\left(\frac{d^{k}}{d^{\underline{k}}}\right),$ (171)
where $\mathbf{X}\sim r(\mathbf{x})$.